OpenStudy (anonymous):

find k such that f(x)=x^3-kx^2+kx+6 has the factor x-2

myininaya (myininaya):

If x-2 is a factor of f then f(2)=0

myininaya (myininaya):

So use this f(2)=0 to find k

OpenStudy (anonymous):

what the question is asking is to find the value of k

myininaya (myininaya):

So you got it @Sona ?

OpenStudy (anonymous):

thanks

myininaya (myininaya):

If you want to put your answer, I will check it

OpenStudy (anonymous):

i'm stuck

OpenStudy (anonymous):

so far i'm @ -2k^2+2k+14=0

myininaya (myininaya):

\[f(2)=0=2^3-k(2)^2+k(2)+6\]

myininaya (myininaya):

hmmm... I don't see any k squared in your problem \[0=8-k(4)+2k+6\] \[0=8-4k+2k+6\]

myininaya (myininaya):

\[0=8-2k+6 \]

myininaya (myininaya):

You can finish the rest I believe right?

OpenStudy (anonymous):

so k shud equal 7?

OpenStudy (anonymous):

ya thats correct thank you!!!

myininaya (myininaya):

yes lol that is right I agree with your homework checker